A vertical line drawn through a normal distribution at z = –1.00 separates the distribution into two sections, the body and the tail. What proportion of the distribution is in the body? Group of answer choices 0.3413 0.1587 0.8413 -0.1587
Answer #1

Solution :

Using standard normal table,

P(z > -1.00) = 1 - P(z < -1.00) = 1 - 0.1587 = 0.8413

Proportion = 0.8413